# The Beilinson–Bloch conjecture for some non-isotrivial varieties over global function fields

### v2

@@ 5. Monodromy of Lefschetz pencils
Here we compile some needed results from [SGA7-II] and [Deligne1980] on the monodromy of Lefschetz pencils, and use them to compute the Hasse–Weil $L$-function arising from the middle cohomology of a Lefschetz pencil over a finite field. This will later shed light on the arithmetic of the intermediate Jacobians of certain cubic threefolds (proposition 7.1).
Let $k$ be a field and $X$ a smooth, projective, geometrically connected variety over $k$ of dimension $d\geq 1$. We fix a closed immersion $X\subset\mathbb{P}^{r}_{k}$ [^5] In concrete examples, the specific closed immersion we consider will typically be clear from context.. Let $\check{\mathbb{P}}^{r}_{k}$ denote the dual projective space parametrizing hyperplanes in $\mathbb{P}^{r}_{k}$, and let $\operatorname{Gr}(1,\check{\mathbb{P}}^{r}_{k})$ denote its Grassmannian of lines. A line $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(k)$ has an associated axis $A(h)=\bigcap\limits_{t\in\mathbb{P}^{1}(k)}h(t)$, which is a codimension-2 linear subspace in $\mathbb{P}^{r}_{k}$.
Definition 5.1 ([SGA7-II, XVII.2.2]).
@@
    for all $t\in\mathbb{P}^{1}(\overline{k})-U(\overline{k})$, $h(t)$ intersects $X_{\overline{k}}$ transversely except at a single point, which is an ordinary double point singularity of $X_{\overline{k}}\cap h(t)$ [SGA7-II, Définition XV.1.2.1]. We also refer to such a singularity as a node.
@@
An ordinary double point singularity of a variety $Z$ over $k$ is non-degenerate in the sense of [SGA7-II, Définition XV.1.2.2] iff either $\dim Z$ is odd, or $\operatorname{char}(k)\neq 2$. Several of the results we will state have similar hypotheses on dimension and characteristic, which are partly explained by the need to address the edge case of degenerate ordinary double points.
We proceed to recall a criterion for the existence of Lefschetz pencils (over $\overline{k}$). Let $N$ be the normal bundle of $X$ in the ambient projective space. Then the associated projective bundle $\mathbb{P}(N)$ embeds into $X\times\check{\mathbb{P}}^{r}_{k}$, and its image consists of pairs $(x,H)$, with $H$ a hyperplane tangent to $X$ at $x$. The image of the projection $\phi:\mathbb{P}(N)\to\check{\mathbb{P}}^{r}_{k}$ is called the dual variety of $X$, denoted by $\check{X}$ [SGA7-II, XVII.3.1.3].
Lemma 5.3 ([SGA7-II, Corollaire XVII.3.5.0]).
Assume that $d$ is even, or that $\operatorname{char}(k)\neq 2$. Then $X_{\overline{k}}\subset\mathbb{P}^{r}_{\overline{k}}$ is a Lefschetz embedding if and only if either $\dim\check{X}<r-2$, or the extension of function fields $k(\mathbb{P}(N))/k(\check{X})$ is finite separable.
When $X$ is a smooth cubic fourfold and $\operatorname{char}(k)\notin\{2,3\}$, the conditions of the lemma are satisfied [CP2019, Lemme 3.2].
Fix a Lefschetz pencil $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(k)$ for $X$, and let $\Delta=\Delta(h)=X\cap A(h)$ be the base locus of the pencil. The blowup $Y$ of $X$ along $\Delta$ is smooth, and admits a morphism $f:Y\to\mathbb{P}^{1}$ such that $f^{-1}(t)\cong X_{\overline{k}}\cap h(t)$ for all $t\in\mathbb{P}^{1}(\overline{k})$ [SGA7-II, XVIII.3.1]. We will slightly abuse terminology and also refer to $f$ as a Lefschetz pencil for $X$. Let $\eta$ be the generic point of $\mathbb{P}^{1}_{k}$, and $\overline{\eta}\to\eta$ a geometric generic point. Let $G_{\eta}=\operatorname{Gal}(\overline{\eta}/\eta)$, and let $V=H^{d-1}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell})$.
@@
This follows from [SGA7-II, Théorèmes XVIII.5.6 and XVIII.6.3.3] and the hard Lefchetz theorem [Deligne1980, Théorème 4.1.1]. ∎
@@
Let $G_{\infty}$ be the Galois group of $\overline{k(\eta)}$ over $\overline{k}(\eta)$, and note that the $G_{\infty}$-action on $V$ factors through the surjection $G_{\infty}\to\pi_{1}(U,\overline{\eta})$. Let $\rho:G_{\infty}\to\mathcal{G}$ denote the $G_{\infty}$-representation on $\operatorname{Ev}(Y_{\overline{\eta}})$. This lands in the group $\mathcal{G}$ of automorphisms of $\operatorname{Ev}(Y_{\overline{\eta}})$ which preserve the nondegenerate cup product pairing ${\operatorname{Ev}(Y_{\overline{\eta}})\otimes_{{\mathbb{Q}}_{\ell}}\operatorname{Ev}(Y_{\overline{\eta}})}\to{\mathbb{Q}}_{\ell}(-(d-1))$.
Proposition 5.6 ([Deligne1980, Théorème 4.4.1]).
@@
Fix $\overline{x}\in S$, and a path $c$ from $\overline{\eta}$ to $\overline{\eta}_{\overline{x}}$. Let $I_{c}\subset G_{\eta}$ be the associated inertia group. Recall the specialization map $\operatorname{sp}:H^{d-1}(Y_{\overline{x}},{\mathbb{Q}}_{\ell})\to H^{d-1}(Y_{\overline{\eta}_{\overline{x}}},{\mathbb{Q}}_{\ell})$ [stacks-project, Tag 0GJ2], whose target is identified via $c$ with $V$.
@@
The claims are immediate from [SGA7-II, Théorème XV.3.4]. ∎
@@
$L(V,s)=\prod\limits_{x\in(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})^{0}}\chi(V^{I_{x}},s)$
@@
$\Lambda(s)=\prod\limits_{x\in(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})^{0}}\chi(H^{d-1}(Y_{\overline{x}},{\mathbb{Q}}_{\ell}),s)$
@@
so that the Euler factors of $L(V,s)$ and $\Lambda(s)$ at each closed point agree. By the Grothendieck–Lefschetz trace formula [Milne1980, Theorem VI.13.3],
$\Lambda(s)=\frac{\chi(H^{1}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)}{\chi(H^{0}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)\chi(H^{2}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)}.$ (5.1)
@@ 6. Lefschetz pencils of some special varieties
Let $k={\mathbb{F}}_{q}$, and again let $X\subset\mathbb{P}^{r}_{k}$ be a smooth, projective, geometrically connected variety of dimension $d$. Fix a Lefschetz pencil $h$ for $X$ which is defined over $k$. Again let $\Delta$ be the base locus of $h$, let $f:Y\to\mathbb{P}^{1}_{k}$ be the family of hyperplane sections of $X$ associated with $h$, and let $Y_{\eta}$ be the generic fiber of $f$. Our goal is to use the criterion of proposition 4.13 to prove cases of the Beilinson–Bloch conjecture for $Y_{\eta}$. This requires understanding the Chow groups of $Y_{\eta}$, and those of the singular closed fibers of $f$. We will give two similar arguments along these lines, first when $X$ is a cubic fourfold, then when $X$ is a more general cubic hypersurface. When $X$ is a threefold with $\operatorname{CH}^{2}(X)$ one-dimensional, we prove that the group of zero-cycles of degree zero of $Y_{\eta}$ is torsion through more elementary means.
@@
$\operatorname{CH}_{d-j}(X_{Z})\to H_{2d-2j}((X_{Z})_{\overline{k}},{\mathbb{Q}}_{\ell}(d-j))^{\operatorname{Frob}_{q}\sim 1}$
@@
$\operatorname{CH}_{d-j}((X_{Z})_{L})\to H_{2d-2j}((X_{Z})_{\overline{k}},{\mathbb{Q}}_{\ell}(d-j))^{\operatorname{Frob}_{q}^{r}\sim 1}$
is surjective. Fix such an $L$ so that the points of $Z$ are $L$-rational. Then $(X_{Z})_{L}$ is a finite disjoint union of hyperplane sections of $X_{L}$, and the desired surjectivity follows from the weak Lefschetz theorem (proposition 4.4). ∎
@@
Suppose that $X$ is a smooth cubic hypersurface of dimension $d\geq 5$. Then $\operatorname{TB}^{4,2}$ holds for all closed fibers of $f$. If $\operatorname{TB}^{6,3}(X)$ holds (e.g. $X$ is a Fermat cubic hypersurface satisfying the hypotheses of corollary 4.8), then $BB^{3}(Y_{\eta})$ does as well.
@@
If $F$ is smooth, then $\operatorname{TB}^{4,2}(F)$ holds by proposition 4.10. We may thus assume that $F$ is smooth away from a single $k$-rational ordinary double point. With $\alpha:B\to F$ the blowup of the singular point, we see from proposition 3.3, lemma 6.3, and the blowup formula for motives that $\operatorname{TB}^{4,2}(B)$ holds. In detail, we have
@@
where $S\subset\mathbb{P}^{d-1}$ is a smooth complete intersection of type (3, 2). Thus $\operatorname{TB}^{2,1}(S)$ implies $\operatorname{TB}^{4,2}(B)$. When $d=5$, it is well-known that $S$ is a K3 surface [Huybrechts2016, Example 1.1.3], so that $\operatorname{TB}^{2,1}(S)$ follows from the known Tate conjecture for $S$ [KM2016, Theorem A.1]. When $d>5$, we observe from the Chow–Künneth decomposition of proposition 3.3 that $\operatorname{TB}^{2,1}(S)$ holds.
@@
As $Y$ is the blowup of $X$ along a smooth curve, the blowup formula for motives shows that $\operatorname{CH}^{2}(Y)$ is two-dimensional. Let $D,L\in\operatorname{CH}^{1}(Y)$, such that $D$ is the class of a closed fiber of $f$, and $L$ is ample. Then $D\cdot L$ is numerically nontrivial in $Y$. Since $D\cdot L$ is in the kernel of the surjection $\operatorname{CH}^{2}(Y)\to\operatorname{CH}^{2}(Y_{\eta})$, its target (which is nonzero) must be one-dimensional. ∎
@@
The statement of proposition 6.8 was previously used in [Kahn2021, Proof of Corollary 1] to prove finiteness of the Albanese kernel of some isotrivial surfaces over global function fields.
@@
Then $Y_{\eta}$ is not isotrivial. In the former case, this follows from propositions 5.5 and 5.6. In the latter case, it follows from [OV2007, Example 9]. Taking $X$ to be the Fermat quartic threefold in corollary 6.9, we thus find that for $h$ sufficiently generic, $Y_{\eta}$ is a non-isotrivial quartic K3 surface over ${\mathbb{F}}_{q}(t)$ whose group of zero-cycles of degree zero is finite. As remarked in [Huybrechts2014, Example 9.3], it would be interesting to find an example of a non-isotrivial K3 surface $Z$ over $\overline{{\mathbb{F}}_{q}(t)}$ with $\operatorname{CH}^{2}(Z)_{0}$ trivial.
6.1. Motivic cohomology groups of $Y_{\eta}$
In this subsection, we assume that $X$ is a smooth cubic fourfold over $k={\mathbb{F}}_{q}$ satisfying $\operatorname{TB}(X)$. Using theorem 6.4 and results of Jannsen, we can compute all motivic cohomology groups of $Y_{\eta}$ (with rational coefficients) in terms of certain simpler invariants, whose definitions we now recall.
Let $K=k(\eta)$. For any smooth variety $W$ over $K$ and $b,$ $c\in{\mathbb{Z}}$, define the arithmetic étale cohomology

### v3

@@ 5. Monodromy of Lefschetz pencils
Here we compile some needed results from [27] and [29] on the monodromy of Lefschetz pencils, and use them to compute the Hasse–Weil $L$-function arising from the middle cohomology of a Lefschetz pencil over a finite field. This will later shed light on the arithmetic of the intermediate Jacobians of certain cubic threefolds (proposition 7.1). The content of this section is applied primarily in section 7.
Let $k$ be a field and $X$ a smooth, projective, geometrically connected variety over $k$ of dimension $d\geq 1$. We fix a closed immersion $X\subset\mathbb{P}^{r}_{k}$ [^6] In concrete examples, the specific closed immersion we consider will typically be clear from context.. Let $\check{\mathbb{P}}^{r}_{k}$ denote the dual projective space parametrizing hyperplanes in $\mathbb{P}^{r}_{k}$, and let $\operatorname{Gr}(1,\check{\mathbb{P}}^{r}_{k})$ denote its Grassmannian of lines. A line $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(k)$ has an associated axis $A(h)=\bigcap\limits_{H\in h}H$, which is a codimension-2 linear subspace in $\mathbb{P}^{r}_{k}$.
Definition 5.1 ([27, XVII.2.2]).
@@
    for all $t\in\mathbb{P}^{1}(\overline{k})-U(\overline{k})$, $h(t)$ intersects $X_{\overline{k}}$ transversely except at a single point, which is an ordinary double point singularity of $X_{\overline{k}}\cap h(t)$ [27, Définition XV.1.2.1]. We also refer to such a singularity as a node.
@@
An ordinary double point singularity of a variety $Z$ over $k$ is non-degenerate in the sense of [27, Définition XV.1.2.2] iff either $\dim Z$ is odd, or $\operatorname{char}(k)\neq 2$. Several of the results we will state have similar hypotheses on dimension and characteristic, which are partly explained by the need to address the edge case of degenerate ordinary double points.
We proceed to recall a criterion for the existence of Lefschetz pencils (over $\overline{k}$). Let $N$ be the normal bundle of $X$ in the ambient projective space. Then the associated projective bundle [^7] Note that [27, Exposé XVII] uses the dual convention for projective bundles, i.e. for a vector bundle $V$, it defines $\mathbb{P}(V)$ to be the bundle of lines in $V^{\vee}$. We use the convention of [95, Tag 01OA]. $\mathbb{P}(N)$ embeds into $X\times\check{\mathbb{P}}^{r}_{k}$, and its image consists of pairs $(x,H)$, with $H$ a hyperplane tangent to $X$ at $x$. The image of the projection $\phi:\mathbb{P}(N)\to\check{\mathbb{P}}^{r}_{k}$ is called the dual variety of $X$, denoted by $\check{X}$ [27, XVII.3.1.3].
Lemma 5.3 ([27, Corollaire XVII.3.5.0]).
Assume that $d$ is even, or that $\operatorname{char}(k)\neq 2$. Then $X_{\overline{k}}\subset\mathbb{P}^{r}_{\overline{k}}$ is a Lefschetz embedding if and only if either $\dim\check{X}\leq r-2$, or the extension of function fields $k(\mathbb{P}(N))/k(\check{X})$ is finite separable.
Corollary 5.4.
Assume that $d$ is even, or that $\operatorname{char}(k)\neq 2$. If $X$ is a smooth hypersurface of degree $e$ in $\mathbb{P}^{r}_{k}$, and $\operatorname{char}(k)\nmid e(e-1)$, then $X_{\overline{k}}\subset\mathbb{P}^{r}_{\overline{k}}$ is a Lefschetz embedding.
Proof.
Since $X$ is a smooth hypersurface, the normal bundle $N$ of $X$ has rank one, and $\mathbb{P}(N)\cong X$. By lemma 5.3, we may assume that $\dim\check{X}=r-1=\dim X$. Then the extension of function fields $k(X)/k(\check{X})$ induced by $\phi$ is finite. Again by the lemma, it suffices to show that $p=\operatorname{char}(k)$ does not divide the degree $k(X):k(\check{X})]$, which we denote by $\deg(\phi)$.
Since $\dim\check{X}=r-1$, $\check{X}$ is a hypersurface in $\check{\mathbb{P}}^{r}_{k}$. Katz in [27, Formula (XVII.5.2.8)] gives the following formula for its degree: we have
$\deg(\check{X})\deg(\phi)=\deg((\xi-[H])^{r-1}),$
where $\xi$ is the divisor of the normal bundle $N$ on $X$, $H$ is a hyperplane section of $X$, and $\deg((\xi-[H])^{r-1})$ denotes the degree of a zero-cycle on $X$. Because $X$ is a smooth hypersurface of degree $e$ in $\mathbb{P}^{r}_{k}$, we have $\xi=e[H]$, and
$\deg(\check{X})\deg(\phi)=\deg(((e-1)[H])^{r-1})=(e-1)^{r-1}\deg([H]^{r-1})=(e-1)^{r-1}e.$
By our hypothesis that $p\nmid e(e-1)$, we see that $p\nmid\deg(\phi)$, as desired. ∎
Fix a Lefschetz pencil $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(k)$ for $X$, and let $\Delta=\Delta(h)=X\cap A(h)$ be the base locus of the pencil. The blowup $Y$ of $X$ along $\Delta$ is smooth, and admits a morphism $f:Y\to\mathbb{P}^{1}$ such that $f^{-1}(t)\cong X_{\overline{k}}\cap h(t)$ for all $t\in\mathbb{P}^{1}(\overline{k})$ [27, XVIII.3.1]. We will slightly abuse terminology and also refer to $f$ as a Lefschetz pencil for $X$. Let $\eta$ be the generic point of $\mathbb{P}^{1}_{k}$, and $\overline{\eta}\to\eta$ a geometric generic point. Let $G_{\eta}=\operatorname{Gal}(\overline{\eta}/\eta)$, and let $V=H^{d-1}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell})$.
@@
This follows from [27, Théorèmes XVIII.5.6 and XVIII.6.3.3] and the hard Lefschetz theorem [29, Théorème 4.1.1]. ∎
@@
Let $G_{\infty}$ be the Galois group of $\overline{k(\eta)}$ over $\overline{k}(\eta)$, and note that the $G_{\infty}$-action on $V$ factors through the map $G_{\infty}\to\pi_{1}(U,\overline{\eta})$, which is surjective by [95, Tag 0BQI]. Let $\rho:G_{\infty}\to\mathcal{G}$ denote the $G_{\infty}$-representation on $\operatorname{Ev}(Y_{\overline{\eta}})$. This lands in the group $\mathcal{G}$ of automorphisms of $\operatorname{Ev}(Y_{\overline{\eta}})$ which preserve the nondegenerate cup product pairing ${\operatorname{Ev}(Y_{\overline{\eta}})\otimes_{{\mathbb{Q}}_{\ell}}\operatorname{Ev}(Y_{\overline{\eta}})}\to{\mathbb{Q}}_{\ell}(-(d-1))$.
Proposition 5.7 ([29, Théorème 4.4.1]).
@@
Fix $\overline{x}\in S$, and a path $c$ from $\overline{\eta}$ to $\overline{\eta}_{\overline{x}}$. Let $I_{c}\subset G_{\eta}$ be the associated inertia group. Recall the specialization map $\operatorname{sp}:H^{d-1}(Y_{\overline{x}},{\mathbb{Q}}_{\ell})\to H^{d-1}(Y_{\overline{\eta}_{\overline{x}}},{\mathbb{Q}}_{\ell})$ [95, Tag 0GJ2], whose target is identified via $c$ with $V$.
@@
The claims are immediate from [27, Théorème XV.3.4]. ∎
@@
$L(V,s)=\prod\limits_{x\in(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})^{0}}\chi(V^{I_{x}},s)^{-1}$
@@
$\Lambda(s)=\prod\limits_{x\in(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})^{0}}\chi(H^{d-1}(Y_{\overline{x}},{\mathbb{Q}}_{\ell}),s)^{-1}$
@@
so that the Euler factors of $L(V,s)$ and $\Lambda(s)$ at each closed point agree. By the Grothendieck–Lefschetz trace formula [77, Theorem VI.13.3],
(5.1) $\Lambda(s)=\frac{\chi(H^{1}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)}{\chi(H^{0}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)\chi(H^{2}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)}.$
@@ 6. Lefschetz pencils of some special varieties
Let $k={\mathbb{F}}_{q}$, and again let $X\subset\mathbb{P}^{r}_{k}$ be a smooth, projective, geometrically connected variety of dimension $d$. Fix a Lefschetz pencil $h$ for $X$ which is defined over $k$. Again let $\Delta$ be the base locus of $h$, let $f:Y\to\mathbb{P}^{1}_{k}$ be the family of hyperplane sections of $X$ associated with $h$, and let $Y_{\eta}$ be the generic fiber of $f$. Our goal is to use the criterion of proposition 4.14 to prove cases of the Beilinson–Bloch conjecture for $Y_{\eta}$. This requires understanding the Chow groups of $Y$, and those of the singular closed fibers of $f$. We will give two similar arguments along these lines, first when $X$ is a cubic fourfold, then when $X$ is a more general cubic hypersurface. When $X$ is a threefold with $\operatorname{CH}^{2}(X)$ one-dimensional, we prove that the group of zero-cycles of degree zero of $Y_{\eta}$ is torsion through more elementary means.
@@
$\operatorname{CH}_{d-j}(Y_{Z})\to H_{2d-2j}((Y_{Z})_{\overline{k}},{\mathbb{Q}}_{\ell}(d-j))^{\operatorname{Frob}_{q}\sim 1}$
@@
$\operatorname{CH}_{d-j}((Y_{Z})_{L})\to H_{2d-2j}((Y_{Z})_{\overline{k}},{\mathbb{Q}}_{\ell}(d-j))^{\operatorname{Frob}_{q}^{r}\sim 1}$
is surjective. Fix such an $L$ so that the points of $Z$ are $L$-rational. Then $(Y_{Z})_{L}$ is a finite disjoint union of hyperplane sections of $X_{L}$, and the desired surjectivity follows from the weak Lefschetz theorem (proposition 4.4). ∎
@@
With this, we can prove the below theorem, which implies theorem 1.1.
@@
Suppose that $X$ is a smooth cubic hypersurface of dimension $d\geq 5$. Then $\operatorname{TB}^{4,2}$ holds for all closed fibers of $f$. If $\operatorname{TB}^{6,3}(X)$ holds (e.g. $X$ is a Fermat cubic hypersurface satisfying the hypotheses of corollary 4.8), then $\operatorname{BB}^{3}(Y_{\eta})$ does as well.
@@
If $F$ is smooth, then $\operatorname{TB}^{4,2}(F)$ holds by proposition 4.10. We may thus assume that $F$ is smooth away from a single $k$-rational ordinary double point. With $\alpha:B\to F$ the blowup of the singular point, we see from proposition 3.7 and lemma 6.3, and the blowup formula for motives, that $\operatorname{TB}^{4,2}(B)$ holds. In detail, we have
@@
where $S\subset\mathbb{P}^{d-1}$ is a smooth complete intersection of type (3, 2). Thus $\operatorname{TB}^{2,1}(S)$ implies $\operatorname{TB}^{4,2}(B)$. When $d=5$, it is well-known that $S$ is a K3 surface [45, Example 1.1.3], so that $\operatorname{TB}^{2,1}(S)$ follows from the known Tate conjecture for $S$ [59, Theorem A.1] and proposition 4.5(iv). When $d>5$, we observe from the Chow–Künneth decomposition of proposition 3.7 that $\operatorname{TB}^{2,1}(S)$ holds.
@@
As $Y$ is the blowup of $X$ along a smooth curve, the blowup formula for motives shows that $\operatorname{CH}^{2}(Y)$ is two-dimensional. Let $L,M\in\operatorname{CH}^{1}(Y)$, such that $M$ is the class of a closed fiber of $f$, and $L$ is ample. Then $L\cdot M$ is numerically nontrivial in $Y$. Since $L\cdot M$ is in the kernel of the surjection $\operatorname{CH}^{2}(Y)\to\operatorname{CH}^{2}(Y_{\eta})$, its target (which is nonzero) must be one-dimensional. ∎
@@
The statement of proposition 4.12 was previously used in [51, Proof of Corollary 1] to prove finiteness of the Albanese kernel of some isotrivial surfaces over global function fields.
@@
Then $Y_{\eta}$ is not isotrivial. In the former case, this follows from propositions 5.6 and 5.7. In the latter case, it follows from [97, Example 9]. Taking $X$ to be the Fermat quartic threefold in corollary 6.9, we thus find that for $h$ sufficiently generic, $Y_{\eta}$ is a non-isotrivial quartic K3 surface over ${\mathbb{F}}_{q}(t)$ whose group of zero-cycles of degree zero is finite. As remarked in [44, Example 9.3], it would be interesting to find an example of a non-isotrivial K3 surface $Z$ over $\overline{{\mathbb{F}}_{q}(t)}$ with $\operatorname{CH}^{2}(Z)_{0}$ trivial.
6.1. Motivic cohomology of $Y_{\eta}$
In this subsection, we assume that $X$ is a smooth cubic fourfold over $k={\mathbb{F}}_{q}$ satisfying $\operatorname{TB}(X)$: for instance, we may take $X$ as in corollary 4.11. Using theorem 6.4 and results of Jannsen, we can compute all motivic cohomology groups of $Y_{\eta}$ (with rational coefficients) in terms of certain simpler invariants, whose definitions we now recall.
Fix a geometric point $\overline{\eta}\to\eta$, and let $K=k(\eta)$. For any smooth variety $W$ over $k(\eta)$ and $b,$ $c\in{\mathbb{Z}}$, define the arithmetic étale cohomology